OpenStudy (anonymous):

2|2x-3|+12 greater than 18

OpenStudy (anonymous):

\[2\left| 2x-3 \right|+12>18\] \[2\left| 2x-3 \right|>6\] \[\left| 2x-3 \right|>3\] the you split to get 2x - 3 > 3 or 2x-3<-3 x > 3 or x<0

OpenStudy (anonymous):

thanks